OpenStudy (anonymous):

Sara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 390 customers ordered turkey, 345 customers ordered ham, and 300 customers ordered neither turkey nor ham. Based on this data, how many of the next 1,000 customers will order both turkey and ham? Show your work, and use complete sentences.

OpenStudy (anonymous):

@vera_ewing

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@sammixbᴏᴏ

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

I don't even know how to go about this. Do I just add 390 and 345? Or is it more complicated than that?

OpenStudy (anonymous):

okay stumped?

OpenStudy (anonymous):

I am

OpenStudy (anonymous):

okay - 1,000 with 390.

OpenStudy (anonymous):

subtract them?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

610

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

ok hold do u have answer choices??

OpenStudy (anonymous):

no this is short response :/

OpenStudy (anonymous):

oh. Ummm hold on will i figure it out ok?

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

where the x represents the number of customers out of 1000 that ordered turkey and ham.

OpenStudy (anonymous):

390 - 300 = 90 people ordered just turkey (and not ham)

OpenStudy (anonymous):

"345 customers ordered ham" so 345 - 300 = 45 just ordered ham (and not turkey next 90+45 = 135 orders were for either turkey or ham (picking one, not both) 300-135 = 165 is for both at the same time so 165

OpenStudy (anonymous):

35 i kinda shorten it but heirs the quick way with calculator. 1000-390=610 345 - 610= 265 300- 265=35 and 35 is the answer.

OpenStudy (anonymous):

so have any more???

OpenStudy (anonymous):

I'm trying to understand this but I see where you're going with it

OpenStudy (anonymous):

so that means u have it?/?

OpenStudy (anonymous):

did u typ down the answer?

OpenStudy (anonymous):

OH okay wait so it would be the same as adding 390, 345, and 300 then subtracting 1000 from 1035

OpenStudy (anonymous):

yes i have it now

OpenStudy (anonymous):

yea every good thinking.

OpenStudy (anonymous):

okay can you help with another? i will make a new post

OpenStudy (anonymous):

I`ll give u a medal for that.

OpenStudy (anonymous):

oh yea.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome.